OpenStudy (candy13106):

which expression is equivalent to (-2) to the 8th power

OpenStudy (anonymous):

(-2)^8=2^8=256

OpenStudy (candy13106):

yea wat is equivalent to it a) (-2)^2 (-2)^4 b) (-2) (-2) ^7 c) [(-2)^4]^4 d) [(-2)^5]^3

OpenStudy (here_to_help15):

can ya draw it out

OpenStudy (candy13106):

ok like how

OpenStudy (anonymous):

(-2) (-2) ^7 addexponets you get 8

OpenStudy (candy13106):

oh ok so its b right

OpenStudy (anonymous):

yeah

OpenStudy (candy13106):

k thanks